# LDI: Localized Data Imputation

### v1

@@ V Performance Analysis
We build a Generalized Suffix Tree (GST) for the $n$ input strings, each of length $\ell$, after appending a unique terminal symbol to each string. This tree can be constructed in $\mathcal{O}(n\cdot\ell)$ time and space using Ukkonen’s algorithm [45]. Each leaf is labeled with its corresponding string ID, and a post-order traversal is used to annotate each internal node with the set of string IDs in its subtree. For each node, if the number of distinct string IDs is at least $\lceil qn\rceil$, we report the substring represented by that node. Reporting all such substrings takes $\mathcal{O}(r)$ time, where $r$ is the number of substrings found. Therefore, the total time and space complexity is $\mathcal{O}(n\cdot\ell+r)$ and $\mathcal{O}(n\cdot\ell)$, respectively. ∎
@@
V-A Attribute Selection
The time complexity of the attribute selection phase depends on the number of candidate attributes ($a_{c}$), the number of groups formed by values in the target attribute ($m$), the number of sampled tuples in each group ($n$), and the average string length in attribute values ($\ell$). The algorithm consists of four main steps: (1) group sampling, (2) pattern detection, (3) uniqueness checking across groups, and (4) an optional redundancy filtering step.
@@ V-C Imputation Using LLM
VI Experiments
In this section, we detail an extensive set of experiments designed to evaluate our proposed data imputation method. Our goal is to demonstrate how selecting relevant attributes and tuples affects the performance and explainability of the imputation process.
VI-A Datasets
Our experiments were conducted on four real-world datasets from three different domains, as summarized in Table II. These datasets vary in size, number of attributes, and incomplete data characteristics, reflecting real-world diversity. The incomplete attributes always contain categorical values, with vocabulary sizes ranging from 49 in the Restaurant dataset to 385 in the Phone dataset. Buy and Restaurant contain fewer attributes compared to our other datasets but are widely used in data imputation studies[13, 46, 8, 33], making them valuable for comparison purposes. In contrast, Zomato and Phone datasets include a larger number of attributes, enabling a more comprehensive evaluation of attribute selection.
TABLE II: Dataset summary
@@ VI-A Datasets
VI-B Experimental Setup
The experiments were conducted on a machine with the following specifications: Ubuntu 20.04, AMD Ryzen 9 3900X 12-Core Processor, and 64 GB of memory, using Python version 3.9.20. For the Buy and Restaurant datasets, all available rows were utilized, and for other datasets, 1000 rows were randomly selected to manage the computational cost and latency of LLM calls. To ensure a fair comparison, the same set of randomly sampled rows was used across different settings and each experiment was repeated and averaged at least five times. In the group sampling phase, we set $m=10$ and $n=10$, ensuring $m$ groups that each contained $n$ samples. In cases where the dataset had fewer groups or samples ($<10$), the maximum available was used. The inner threshold $q$ and outer threshold $p$ were both set to 0.9 for the Buy and Phone datasets, and to 0.8 for the Zomato and Restaurant datasets, as these contain more inconsistencies and variation in their values. This setup allows for slight flexibility in pattern matching—e.g., with a 0.9 threshold, up to 1 row per group and up to 1 group across all groups may have no patterns—while still enforcing strong evidence of dependency. For the LLM-based imputation, we employed GPT-4o-mini (version 2024-07-18) for all experiments, and also included Llama 3.2 3B to evaluate performance in limited-resource settings. While GPT-4o-mini is a larger, high-performing model, Llama 3.2 3B is an open-source alternative with fewer parameters (3 billion) and reduced computational requirements. By including Llama in our experiments, we aim to demonstrate the effectiveness and generalizability of our method even when deployed with smaller, more accessible language models.
VI-C Data Reduction
Table III highlights the efficiency of our model in reducing data size by focusing only on relevant attributes. Unlike other methods that use all available attributes, such as FMW, our approach selects a smaller set of dependent attributes identified in the initial phase. This targeted selection leads to a significant reduction in the amount of data processed, as shown by the lower character counts across all datasets. The reduction in characters is computed by comparing the average length of the selected attributes to that of all attributes. The effect is substantial across all benchmarks: in the Zomato dataset, character input is reduced by 95.7% through the selection of only 2 out of 16 attributes. Similarly, the Buy, Restaurant, and Phone achieve reductions of 57.1%, 39.9%, and 74.0%, respectively. As a result, the process becomes more explainable and the outcomes more traceable, since it is easier to understand which attributes contribute to the final results.
TABLE III: Average reduction in number of attributes and characters
@@ VI-C Data Reduction
VI-D Performance Compared to Baselines
We compare LDI with baseline methods (discussed in §$\S$§II) including KNNE [27] and MIBOS [28] (nearest-neighbor-based), CMI [2] (clustering-based), ERACER [20], Baran [47], and HoloClean [21] (statistics-based), along with two approaches that use language models: IPM [15] and FMW [13]. Table IV presents the accuracy of LDI compared to these baselines under a 10% missing value rate across four datasets. Our method, LDI, achieves the highest accuracy on three datasets (Buy, Restaurant, and Phone) and the second-highest on Zomato. LDI, IPM, and FMW all utilize language models and demonstrate high performance across all datasets. This creates a clear gap between them and traditional methods, which may perform reasonably on a single dataset but fail to generalize. For instance, Baran and HoloClean perform well only on Zomato, with substantially lower results on the other datasets.
IPM treats imputation as a classification task and introduces two variants: IPM-Multi and IPM-Binary [^3] Under the IPM column, we report the better result of these two for each dataset.. IPM-Multi formulates the problem as multiclass classification over the candidate set, and IPM-Binary ranks each candidate using a binary classification model. Both require fine-tuning to capture the relation between context and candidate values, making the model resource-intensive and less flexible. Moreover, IPM performance is sensitive to the quality of training data and the chosen modeling strategy. In practice, large and inconsistent gaps are often observed between IPM-Multi and IPM-Binary across datasets. In contrast, LDI avoids the need for training data by dynamically selecting informative attributes and constructing few-shot prompts using relevant examples. This allows LDI to adapt to new datasets more easily and generate more explainable localized prompts that are better aligned with the LLM’s strengths.
FMW performs imputation by prompting a language model with randomly selected tuples from the dataset. All attributes are included in the prompt, regardless of their relevance, which can introduce noise and distract the model. While FMW avoids training and is simple to apply, it does not distinguish between informative and irrelevant data. In contrast, LDI focuses on localized context and selects only the most relevant attributes of similar tuples, leading to more accurate, traceable, and therefore explainable results across datasets.
LDI is designed to maintain high accuracy regardless of the missing value rate, as long as it is provided with a few well-chosen examples. Like FMW, LDI does not rely on training data and instead uses a few examples to guide the language model. For best results, these examples should be sufficient to detect dependencies and also provide contextual information to frame the imputation task. This behavior is similar to Programming by Examples (PBE) methods, where the model generalizes well from a small, informative set of examples.
TABLE IV: LDI accuracy compared to baselines at 10% missing rate
@@ VI-D Performance Compared to Baselines
TABLE V: Impact of attribute selection phase on accuracy with randomly selected tuples
@@
TABLE VI: Impact of the tuple selection phase on accuracy using only the dependent attributes
@@
VI-E Ablation Study
VI-E1 Impact of Attribute Selection Phase
Table V presents results for a variant of LDI in which tuples are randomly sampled, but only attributes that are in $(p,q)$-approximate dependency with the target attribute are retained. This setting allows us to evaluate the standalone effect of removing irrelevant attributes.
VI-E2 Impact of Tuple Selection Phase
Table VI compares two settings: one using randomly selected tuples and the other using the proposed diverse similarity-based method. Note that in our method, the similarity between tuples is explicitly computed over the attributes identified in phase one. This ensures that the selected examples are similar in the most informative dimensions, avoiding irrelevant features that could dilute the closeness of the examples.
@@ VI-F Varying p and q
To study the effect of hyperparameters p and q, we conduct an experiment on the Zomato dataset, which contains 17 attributes. We chose this dataset specifically because the larger number of columns allows us to better observe how the attribute selection step responds to different values of p and q. We vary both parameters from 0 to 0.75 in increments of 0.25, and for each setting, we evaluate the performance of our method over 5 runs. We do not report results for settings where p and q are increased up to 1, as the Zomato dataset is particularly noisy and contains many inconsistencies in its fields. Applying such strict thresholds would result in no attributes being selected in many sampled data, even when some attributes contain informative patterns. When either p or q is set to 0, no constraint is imposed on the the dependency relationship, and all attributes are retained. In particular, setting $q=0$ removes the frequency threshold, allowing any substring that co-occurs with a distinct value in the target column to be accepted as a valid pattern. Similarly, setting $p=0$ eliminates the requirement on the minimum number of distinct target values that must support the dependency.
@@
In Table VII, we evaluate how varying the hyperparameters $p$ and $q$ affects the imputation performance on the Zomato dataset when $k=3$ examples are selected. We compare two example selection strategies: (a) random selection and (b) diverse similarity-based selection. For both strategies, we observe a notable performance gain when $p$ and $q$ are set to 0.25 or higher, compared to settings where either $p$ or $q$ is 0 and all attributes are included. The gains are more pronounced in the random selection setting, where attribute filtering plays a greater role in guiding the model. In the first setting (Table VII(a)), where examples are selected randomly, we observe that using $p=0$ and $q=0$ yields an exact match accuracy of 0.846 and ROUGE-1 score (shown in parentheses) of 0.883. These scores represent the baseline performance without applying our method, and match the results of the FMW. As we increase $p$ and $q$ to 0.25, the performance improves significantly, reaching approximately 0.960 accuracy and 0.978 ROUGE-1. This indicates that even a modest level of filtering can effectively remove distracting features and highly improve model performance. As $p$ and $q$ increase further, performance remains consistently high. Notably, even at $p=0.75$ and $q=0.75$, where the model uses only 2 attributes, reducing nearly 95% of the input content in terms of characters (as discussed next), the accuracy remains close to 0.960. This shows that our method can retain the critical information required for imputation while drastically reducing the input size.
@@
In the second setting (Table VII(b)), where tuples are selected using diverse similarity, the overall performance improves compared to random selection across all values of $p$ and $q$. The accuracy under $p=0$ and $q=0$ reaches 0.944, with a ROUGE-1 score of 0.951, showing that selecting diverse and informative examples enhances performance even without attribute filtering. As $p$ and $q$ increase, performance continues to improve. The best results are achieved at $p=0.75$, $q=0.50$ and $p=0.50$, $q=0.75$, where the model reaches an accuracy of 0.978 and ROUGE-1 scores of 0.990 and 0.986, respectively. However, these settings require more attributes compared to the setting $p=0.75$, $q=0.75$, which uses just 2 attributes (see Table VIII) and achieves a very similar performance. Given that the difference in performance is minimal, the stricter setting with fewer attributes is a more practical choice, as it reduces input size while maintaining nearly identical accuracy. This can highlight the flexibility of our method in allowing users to balance performance with input efficiency.
@@
To study the effect of varying p and q on attribute reduction, we observe that as the values of p and q increase, the algorithm becomes stricter in its selection. The results are shown in Table VIII. When either parameter is set to 0, the method selects all 16 attributes. As both parameters increase, the average number of selected attributes steadily decreases. For instance, at $p=0.50$, $q=0.50$, only 7.6 attributes are selected on average, and this number drops to just 2 when $p=0.75$ and $q=0.75$. This trend confirms that tighter dependency thresholds are effective in filtering out weak or noisy attribute relationships, enabling the model to focus on the most informative signals.
@@
TABLE VII: Effect of varying $(p,q)$ on imputation performance in the Zomato dataset with $k=3$
@@
TABLE VIII: Effect of varying $(p,q)$ on the average # selected Attr. in the Zomato dataset
$\downarrow$   $\rightarrow$ 0 0.25 0.50 0.75
@@
VI-G Limited-Resource Setting

### v2

@@ 5. Performance Analysis
We build a Generalized Suffix Tree (GST) for the $n$ input strings, each of length $\ell$, after appending a unique terminal symbol to each string. This tree can be constructed in $\mathcal{O}(n\cdot\ell)$ time and space using Ukkonen’s algorithm (Ukkonen, 1995). Each leaf is labeled with its corresponding string ID, and a post-order traversal is used to annotate each internal node with the set of string IDs in its subtree. For each node, if the number of distinct string IDs is at least $\lceil qn\rceil$, we report the substring represented by that node. Reporting all such substrings takes $\mathcal{O}(r)$ time, where $r$ is the number of substrings found. Therefore, the total time and space complexity is $\mathcal{O}(n\cdot\ell+r)$ and $\mathcal{O}(n\cdot\ell)$, respectively. ∎
@@
5.1. Attribute Selection
The time complexity of the attribute selection phase depends on the number of candidate attributes ($a_{c}$), the number of groups formed by values in the target attribute ($m$), the number of sampled tuples in each group ($n$), and the average string length in attribute values ($\ell$). The algorithm consists of four steps: (1) an optional group sampling, (2) pattern detection, (3) uniqueness checking across groups, and (4) an optional redundancy filtering step.
@@ 5.3. Imputation Using LLM
6. Experiments
This section presents a comprehensive evaluation of our proposed data imputation framework, LDI. The experiments aim to answer three main questions: (1) How does attribute and tuple selection affect imputation accuracy and explainability? (2) How does LDI compare to existing state-of-the-art baselines? (3) How robust and scalable is LDI across datasets and missingness conditions?
6.1. Datasets
We evaluate LDI on four real-world datasets from three distinct domains, summarized in Table 1. The datasets vary widely in size, number of attributes, and the characteristics of incomplete data, providing a diverse and realistic testbed. The incomplete attributes are categorical, with vocabulary sizes ranging from 49 in the Restaurant dataset to 385 in the Phone dataset. Buy and Restaurant contain fewer attributes compared to our other datasets but are widely used in data imputation studies(Narayan et al., 2022; Chen et al., 2024; Song and Sun, 2020; Mei et al., 2021), making them valuable for comparison purposes. In contrast, Zomato and Phone datasets include a larger number of attributes, enabling a more comprehensive evaluation of attribute selection and scalability.
Table 1. Dataset summary
@@ 6.1. Datasets
6.2. Experimental Setup
All experiments were conducted on an Ubuntu 20.04 machine with an AMD Ryzen 9 3900X 12-Core Processor, and 64 GB of memory, using Python version 3.9.20. For the Buy and Restaurant datasets, all available rows were utilized, while for larger datasets (Zomato, Phone), 1000 rows were randomly selected to manage the computational cost and latency of LLM calls. The same samples were reused across all methods to ensure fairness, and each experiment was repeated five times, with the averaged results reported. Group sampling was configured with $m=10$ groups and $n=10$ samples per group. In cases where the dataset had fewer groups or samples (<10<10), the maximum available was used. Thresholds $p$ and $q$ were both set to 0.9 for Buy and Phone, and 0.8 for Zomato and Restaurant, allowing slight flexibility for noisy or inconsistent datasets. A 0.9 threshold permits up to one unmatched row per group and one group with no pattern, maintaining strong dependency evidence while tolerating minor noise. For LLM-based imputation, we primarily employed GPT-4o-mini (version 2024-07-18), and also included Llama 3.2 3B to evaluate performance in limited-resource settings. While GPT-4o-mini is a larger, high-performing model, Llama 3.2 3B demonstrates LDI remains effective with smaller, open-weight models—highlighting its portability and generalizability.
6.3. Data Reduction via Attribute Selection
Table 2 highlights the efficiency of our model in reducing data size by focusing only on relevant attributes. Unlike other methods that use all available attributes, such as FMW, our approach selects a smaller set of dependent attributes identified in the initial phase. This targeted selection leads to a significant reduction in the amount of data processed, as shown by the lower character counts across all datasets. The reduction in characters is computed by comparing the average length of the selected attributes to that of all attributes. The effect is substantial across all benchmarks: in the Zomato dataset, character input is reduced by 95.7% through the selection of only 2 out of 16 attributes. Similarly, the Buy, Restaurant, and Phone achieve reductions of 57.1%, 39.9%, and 74.0%, respectively. As a result, the process becomes more explainable and the outcomes more traceable, since it is easier to understand which attributes contribute to the final results.
Table 2. Average reduction in number of attributes and characters
@@ 6.3. Data Reduction via Attribute Selection
6.4. Performance Compared to Baselines
We compare LDI with both traditional and LLM-based baselines: KNNE (Domeniconi and Yan, 2004) and MIBOS (Wu et al., 2012) (nearest-neighbor-based), CMI (Zhang et al., 2008) (clustering-based), ERACER (Mayfield et al., 2010), Baran (Mahdavi and Abedjan, 2020), and HoloClean (Rekatsinas et al., 2017) (statistics-based), along with two approaches that use language models: IPM (Li and Döhmen, 2024) and FMW (Narayan et al., 2022). Baran, originally an error correction model, is adapted for data imputation by replacing each missing value with a null token and predicting its correction. The number of labels (examples) is set to 20, consistent with prior studies (Li and Döhmen, 2024). To simulate missing data, we apply a missing completely at random mechanism to originally complete datasets, removing 10% of the values. Table 3 reports the accuracy of LDI compared to these baselines across four datasets. Our method achieves the highest accuracy on three datasets (Buy, Restaurant, and Phone) and the second-highest on Zomato. LDI, IPM, and FMW all utilize language models and demonstrate high performance across all datasets. This creates a clear gap between them and traditional methods, which may perform reasonably on a single dataset but fail to generalize. For instance, Baran and HoloClean perform well only on Zomato, with substantially lower results on the other datasets.
Table 3. LDI accuracy compared to baselines at 10% missing rate
@@ 6.4. Performance Compared to Baselines
IPM treats imputation as a classification task and introduces two variants: IPM-Multi and IPM-Binary [^3] Under the IPM column, we report the better result of these two for each dataset.. IPM-Multi formulates the problem as multiclass classification over the candidate set, and IPM-Binary ranks each candidate using a binary classification model. Both require fine-tuning to capture the relation between context and candidate values, making the model resource-intensive and less flexible. Moreover, IPM performance is sensitive to the quality of training data and the chosen modeling strategy. In practice, large and inconsistent gaps are often observed between IPM-Multi and IPM-Binary across datasets. In contrast, LDI avoids the need for training data by dynamically selecting informative attributes and constructing few-shot prompts using relevant examples. This allows LDI to adapt to new datasets more easily and generate more explainable localized prompts that are better aligned with the LLM’s strengths. FMW performs imputation by prompting a language model with tuples from the dataset, where examples can be randomly or manually selected. All attributes are included in the prompt, regardless of their relevance, which can introduce noise and distract the model. While FMW avoids training, is simple to apply, and allows manual selection of informative examples, this selection does not scale to large datasets [^4] We therefore use random selection for comparison. and their method does not distinguish between informative and irrelevant attributes. In contrast, LDI focuses on localized context and selects only the most relevant attributes of similar tuples, leading to more accurate, traceable, and therefore explainable results across datasets.
6.5. Robustness to Different Missingness Levels
LDI is designed to maintain high accuracy regardless of the missing value rate, as long as it is provided with a few well-chosen examples. For best results, these examples should be sufficient to detect dependencies and also provide contextual information to frame the imputation task. This behavior is similar to Programming by Examples (PBE) methods, where the model generalizes well from a small, informative set of examples. To evaluate the robustness of LDI to changes in missing rate, we vary the missing rate from 10% to 50%, with a missing completely at random mechanism applied to complete tuples. As shown in Figure 4, even after imputing 50% of the data, the accuracy does not drop significantly, indicating that the model can still identify attribute dependencies and use similar examples to guide the LLM. This robustness is expected to extend to different missingness mechanisms—missing at random and missing not at random—because the model can uncover attribute dependencies even from limited samples, as reflected in the results, and a small set of informative examples is sufficient to drive accurate imputations.
Figure 4. Impact of varying missing rates on accuracy
6.6. Runtime and Scalability
Table 4 summarizes the execution time of the attribute and tuple selection phases on our real-world datasets. The attribute selection phase, executed once to identify dependent columns, completes in negligible time across all datasets. The slightly higher values in the Zomato and Phone datasets stem from their larger number of attributes and longer average string lengths. In contrast, the tuple selection phase must be performed for each missing value, and its runtime depends mainly on the number of tuples compared during similarity search, as well as the number and length of attributes selected in the first phase.
To further assess scalability with respect to input characteristics—the number of groups ($m$), tuples per group ($n$), and attribute value length in characters (ll)—we conducted controlled experiments on synthetic data. We varied $l$ from 10 to 1000 and both $m$ and $n$ from 10 to 50. As shown in Section 5, these parameters have the largest impact on the runtime of attribute selection. The number of returned substrings was controlled by setting $p=q=0.5$. The results are shown in Figure 5. As expected, runtime increases with both the number of tuples and the average string length during the attribute selection phase (Figure 5(a)), but remains practical even for lengthy strings and thousands of tuples. Group sampling can further bound this cost by operating on a small yet representative subset of the data; for instance, sampling only 100 tuples ($m=n=10$) was sufficient in our experiments to detect reliable patterns. The runtime of the tuple selection phase exhibits a similar trend (Figure 5(b)), growing moderately as the number and length of tuples increase. Since this phase must be repeated for every missing entry, its cost scales linearly with the number of imputations. When multiple attributes are selected in the first phase, the runtime increases proportionally, as similarity is computed pairwise across attributes. Nevertheless, even in these settings, runtime remains well within practical limits, supporting the scalability of our approach.
Table 4. Runtime on real-world datasets (in seconds)
Dataset Attr. Selection Tup. Selection
Buy 0.022 0.058
Restaurant 0.012 0.116
Zomato 0.338 0.375
Phone 0.795 0.162
(a) Attr. Selection
(b) Tup. Selection
Figure 5. Runtime trends on synthetic data (in seconds)
6.7. Ablation Study
6.7.1. Impact of Attribute Selection Phase
Table 5 presents results for a variant of LDI in which tuples are randomly sampled, but only attributes that are in $(p,q)$-approximate dependency with the target attribute are retained. This setting allows us to evaluate the standalone effect of removing irrelevant attributes.
Table 5. Impact of attribute selection phase on accuracy with randomly selected tuples
@@ 6.7.1. Impact of Attribute Selection Phase
6.7.2. Impact of Tuple Selection Phase
Table 6 compares two settings: one using randomly selected tuples and the other using the proposed diverse similarity-based method. Note that in our method, the similarity between tuples is explicitly computed over the attributes identified in phase one. This ensures that the selected examples are similar in the most informative dimensions, avoiding irrelevant features that could dilute the closeness of the examples.
Table 6. Impact of the tuple selection phase on accuracy using only the dependent attributes
@@ 6.8. Varying p and q
To evaluate the robustness of LDI under varying levels of noise in approximate dependencies, we conducted an experiment where p and q varied from 0 to 0.75 in increments of 0.25, and the performance was evaluated over five independent runs for each setting. The experiment was performed on the Zomato dataset, whose larger number of columns allows us to better observe how the attribute selection step responds to different noise levels. We omit results for settings where p and q exceed 0.75, as the Zomato dataset is particularly noisy and contains many inconsistencies in its fields. Applying stricter thresholds (e.g., p, q =1) would eliminate nearly all dependencies, leaving no attributes selected in many samples, even when some contain informative patterns. When either parameter is set to 0, no constraint is imposed on the dependency relationship, and all attributes are retained. In particular, setting $q=0$ removes the frequency threshold, allowing any substring that co-occurs with a distinct value in the target column to be accepted as a valid pattern. Similarly, setting $p=0$ eliminates the requirement on the minimum number of distinct target values that must support the dependency.
@@
In Table 7, we evaluate how varying the hyperparameters $p$ and $q$ affects the imputation performance on the Zomato dataset when $k=3$ examples are selected. We compare two example selection strategies: (a) random selection and (b) diverse similarity-based selection. For both strategies, we observe a notable performance gain when $p$ and $q$ are set to 0.25 or higher, compared to settings where either $p$ or $q$ is 0 and all attributes are included. The gains are more pronounced in the random selection setting, where attribute filtering plays a greater role in guiding the model. In the first setting (Table 7(a)), where examples are selected randomly, we observe that using $p=0$ and $q=0$ yields an exact match accuracy of 0.846 and ROUGE-1 score (shown in parentheses) of 0.883. These scores represent the baseline performance without applying our method, and match the results of the FMW. As we increase $p$ and $q$ to 0.25, the performance improves significantly, reaching approximately 0.960 accuracy and 0.978 ROUGE-1. This indicates that even a modest level of filtering can effectively remove distracting features and highly improve model performance. As $p$ and $q$ increase further, performance remains consistently high. Notably, even at $p=0.75$ and $q=0.75$, where the model uses only 2 attributes, reducing nearly 95% of the input content in terms of characters (as discussed next), the accuracy remains close to 0.960. This shows that our method can retain the critical information required for imputation while drastically reducing the input size.
@@
In the second setting (Table 7(b)), where tuples are selected using diverse similarity, the overall performance improves compared to random selection across all values of $p$ and $q$. The accuracy under $p=0$ and $q=0$ reaches 0.944, with a ROUGE-1 score of 0.951, showing that selecting diverse and informative examples enhances performance even without attribute filtering. As $p$ and $q$ increase, performance continues to improve. The best results are achieved at $p=0.75$, $q=0.50$ and $p=0.50$, $q=0.75$, where the model reaches an accuracy of 0.978 and ROUGE-1 scores of 0.990 and 0.986, respectively. However, these settings require more attributes compared to the setting $p=0.75$, $q=0.75$, which uses just 2 attributes (see Table 8) and achieves a very similar performance. Given that the difference in performance is minimal, the stricter setting with fewer attributes is a more practical choice, as it reduces input size while maintaining nearly identical accuracy. This can highlight the flexibility of our method in allowing users to balance performance with input efficiency.
@@
To study the effect of varying p and q on attribute reduction, we observe that as the values of p and q increase, the algorithm becomes stricter in its selection. The results are shown in Table 8. When either parameter is set to 0, the method selects all 16 attributes. As both parameters increase, the average number of selected attributes steadily decreases. For instance, at $p=0.50$, $q=0.50$, only 7.6 attributes are selected on average, and this number drops to just 2 when $p=0.75$ and $q=0.75$. This trend confirms that tighter dependency thresholds are effective in filtering out weak or noisy attribute relationships, enabling the model to focus on the most informative signals.
@@
Table 7. Effect of varying $(p,q)$ on imputation performance in the Zomato dataset with $k=3$
@@
Table 8. Effect of varying $(p,q)$ on the average # selected Attr. in the Zomato dataset
[…]
@@
6.9. Limited-Resource Setting